Solve for the variable in the following proportion.
1/4 is to 1 1/4 as 2 is to b
b=

Answer by stanbon(75887)

Solve for the variable in the following proportion.
1/4 is to 1 1/4 as 2 is to b
-------
2/b = (1/4)/(5/4)
-----
Invert:
b/2 = (5/4)/(1/4)
----
b/2 = 5
----
b = 10
-------------
Cheers,
Stan H.
-------------